Lauren always throws loose change into a pencil holder on her desk and takes it out every two weeks. This time it is all nickels and dimes. There are 7 times as many dimes as nickels, and the value of the dimes is $3.90 more than the value of the nickels. How many nickels and dimes does Lauren have?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52782)

.
10*(7x) - 5x = 390   cents


70x - 5x = 390

65x      = 390

  x      = 390/65 = 6.


ANSWER.  6 nickels and 6*7 = 42 dimes.


CHECK.   42*10 - 3*6 = 390 cents.   ! Precisely correct !

Solved.

---------------


Is everything clear to you in my solution ?

If not, let me know.